# HuRef: HUman-REadable Fingerprint for Large Language Models

### v1

@@ 4 Making The Invariant Terms Human-readable: The Fingerprinting Model
4.1 Training the Convolutional Encoder
All the invariant terms in Equation 8 have the same size, i.e., ${\bm{M}}_{a},\displaystyle{\bm{M}}_{b},\displaystyle{\bm{M}}_{f}\in\mathbb{R}^{K\times K}$, regardless of the index of the layer or LLM sizes. As a result, we can tile up them to form a 3D input tensor ${\bm{M}}\in\mathbb{R}^{K\times K\times C}$, where $C$ is the channel dimension. If we are using all layers, $C=3N$. Again, in order to make ${\bm{M}}$ the same size across all models, we only involve the last $r$ layers in the LLM [^1] In fact, experimentally we find that a small $r$ is already sufficient to discriminate LLMs, it’s not necessary to involve many layers. In all of our experiments, $r=2$, so there are only $6$ channels in the input..
Note that we don’t need to use any real LLM weights for training the convolutional encoder, as it only needs to learn a locality-preserving mapping between the input tensor and the output Gaussian vector. This ensures strict exclusivity between the training and test data. To construct the training data, we synthesize the matrix in each channel of ${\bm{M}}$ on-the-fly, by randomly sampling 3 matrices ${\bm{P}}_{1},\displaystyle{\bm{P}}_{2},\displaystyle{\bm{P}}_{3}$ and multiplying them together as ${\bm{P}}_{1}\displaystyle{\bm{P}}_{2}\displaystyle{\bm{P}}_{3}\displaystyle{\bm{P}}_{1}^{T}$, as though they are model parameters.
To learn locality-preserving mapping, we adopt contrastive learning. For a randomly sampled input ${\bm{M}}$, its negative sample is given by another independently sampled tensor ${\bm{M}}^{-}$. For its positive sample ${\bm{M}}^{+}$, we perturb the content in each of ${\bm{M}}$’s channel by adding small perturbation noises $\epsilon_{i}\in\mathcal{N}(0,\alpha)$ to the 3 matrices behind it. Following is the detailed process.
Here $\alpha$ is a hyperparameter determining the small variance. Subsequently, the contrastive loss $\mathcal{L}_{C}$ is given by:
To render the output vector appears to be Gaussian, we adopt the standard GAN (Karras et al., 2019) training scheme. We add a simple MLP as the discriminator $D$ that is trained to discriminate between real Gaussian vectors and the convolutional output vector ${\bm{v}}$. In this setting, the convolutional encoder serves as the generator. During training, for every $m$ steps, we alternate between training the discriminator and the generator. The discriminator loss $\mathcal{L}_{D}$ is thus given by
$\mathcal{L}=\displaystyle\mathcal{L}_{C}+\mathcal{L}_{D}$ (14)
@@ 4.2 Inference
In the inference stage, the convolutional encoder takes the invariant terms from real LLMs and outputs ${\bm{v}}$. One image generator converts ${\bm{v}}$ into a natural image. In principle, any image generator that takes a Gaussian input and has the locality-preserving property would fit here. By visually representing the invariant terms as fingerprints, we can easily identify base models based on their fingerprint images, enabling reliable tracking of model origins. In this paper, we employ the StyleGAN2 generator pretrained on the AFHQ (Choi et al., 2020) dog dataset to generate natural images, following is a detailed description of it.
Figure 5: Fingerprints of 6 different base models (in the first row) and their corresponding offspring models (the lower two rows) are presented. The base model’s name is omitted in the offspring models. The blue number to the right of each offspring model indicates the cosine similarity of its invariant terms (ICS) w.r.t. its base model.
@@ 5 Experiments
Apart from the results shown in Section 3, our experiment section is Four-fold. First, we choose several widely used and open-sourced LLM base models with their offspring LLMs to compute their invariant terms and fingerprint images. Second, we extensively test our method on the LLaMA model family, including some heavily continue-pretrained ones to experimentally verify the robustness of our method against subsequent training processes. Third, we extensively experimented on more open-sourced LLMs to show the diversity of generated images on different LLMs. Finally, we validated the uniqueness and stability of the parameter direction of independently trained LLMs in smaller scale.
5.1 Independently Trained LLMs and Their Offspring Models
We conduct experiments on 6 commonly used open-sourced LLMs with their sizes ranging from 7B to 40B. For the invariant terms, we use the last 2 layers’ parameters and set $K=4096$, which results in ${\bm{M}}\in\mathbb{R}^{4096\times 4096\times 6}$ for all models. The 6 base models are Falcon-40B, MPT-7B (Lin et al., 2022), LLaMA2-7B, Qwen-7B (Bai et al., 2023), Internlm-7B, and Baichuan-13B. Most of their offspring models went through an SFT process, except for LLaMA2-7B-32K (Together.ai, 2023), which also had a continued pretraining before SFT. Following are detailed descriptions of these models.
Falcon-40B-Instruct and Falcon-40B-SFT-Top1-560 underwent SFT on Falcon-40B using data from Baize and top-1 quality demonstrations from the OASST dataset (Köpf et al., 2023) respectively. MPT-7B-Instruct underwent SFT using data from Databricks Dolly-15k (Conover et al., 2023) and the HH-RLHF datasets (Bai et al., 2022), while MPT-7B-StoryWriter was specifically optimized for reading and writing fictional stories. LLaMA2-7B-Chat is the dialogue version of LLaMA2-7B, and LLaMA2-7B-32K underwent continued pretraining and SFT to enhance its long-context capacity. Qwen-7B-Chat was fine-tuned with alignment techniques, and Firefly-Qwen-7B underwent SFT by the Firefly project. Baichuan-13B-Chat is the dialogue version of Baichuan-13B, and Baichuan-13B-SFT is a bilingual instruction-tuned model. InternLM-7B-Chat was optimized for dialogue use cases, and Firefly-InternLM underwent SFT by the Firefly project.
Apart from the fingerprints (Figure 5) we also calculate, for each offspring model, its invariant terms’ cosine similarity (ICS) w.r.t. its base model. Remarkably, for all the offspring models, their fingerprints closely resemble those of their base models, and the ICS indicates the high similarity of their invariant terms w.r.t. their base models. On the other hand, LLMs based on different base models yield fingerprints that are highly different, covering diverse appearances and breeds of dogs.
@@ 5.1 Independently Trained LLMs and Their Offspring Models
Table 3: The cosine similarities of invariant terms between various pairs of LLaMA-7B and its offspring models. Abbreviations used include ”MAlpaca” for ”MedAlpaca,” ”AlpacaL” for ”Alpaca-Lora,” ”MiGPT” for ”MiniGPT-4,” ”Wizard” for ”WizardLM,” and ”CAlpaca” and ”CLLaMA” representing ”Chinese-Alpaca” and ”Chinese-LLaMA,” respectively.
Figure 6: Fingerprints of LLaMA-7B and its offspring models.
5.2 LLaMA and its offspring models
To extensively validate how our proposed method is robust to various subsequent training processes, we choose the LLaMA-7B base model as a testbed, since it is a widely used model with the largest family of offspring models. We collect $10$ offspring models introduced in Section 3.1.1. Plus Beaver which underwent RLHF (Dai et al., 2023), Guanaco (Dettmers et al., 2023) SFT on the multilingual dataset, and BiLLa (Li, 2023) continued pretraining on a new language. Among these offspring models, We want to draw attention to the Chinese-LLaMA and Chinese-Alpaca models, both of which have an extended vocabulary and have been continued pretrained for 20GB of text in a new language, with the latter trained with several millions of additional SFT samples. Following the settings in Section 5.1, the invariant terms again use the last 2 layers’ parameters.
Despite all these complex subsequent training steps, all the models yield highly relevant invariant terms (Table 3). We extensively compute the cosine similarity of the invariant terms (ICS) between every pair of models and found that they are still highly similar, with the lowest ICS being $73.11$. For our trained fingerprinting model, they are aligned to a similar fingerprint image of a Croatian sheepdog, with similar poses, coat patterns, expressions, and backgrounds (Figure 6).
5.3 A wider collection of Independently Trained LLMs
In this experiment, we aimed to gather diverse models covering various parameter sizes. For the widely used LLaMA models, we included LLaMA-7B, LLaMA-13B, LLaMA-65B, and LLaMA2-7B. We also incorporated models with similar architectures to LLaMA, such as InternLM-7B, OpenLLaMA-7B, and Baichuan-7B. To encompass a broader range of parameters, we expanded our collection to include GPT2-Large (Radford et al., 2019), Cerebras-GPT-1.3B (Dey et al., 2023), GPT-Neo-2.7B, and even the largest Falcon-180B. Additionally, we considered models like Qwen-7B, RedPajama-7B (Computer, 2023), ChatGLM-6B (Du et al., 2022), Bloom-7.1B (Workshop et al., 2022), ChatGLM2-6B, Pythia-6.9B and 12B (Biderman et al., 2023), OPT-6.7B and 30B, and GPT-NeoX-20B (Black et al., 2022), among other commonly used LLMs. Finally, We collect a total of 24 open-sourced LLMs with their sizes ranging from 774M (GPT2-Large) to 180B (Falcon-180B). We generate their fingerprints, as shown in Figure 7. The similarities between different models were notably low(Table 4), further confirming the validity of our proposed method and model.
ICS GPT2 CGPT1.3 Neo2.7 CGLM CGLM2 OPT6.7 Pythia6.9 LM7 Qwen LM27 RedPajama Bloom Internlm OLM Baichuan Pythia12 LM13 Neox20 OPT30 LM30 Gala30 LM65 Gala120 Fal180
GPT2 100.00 18.06 -1.07 -0.67 0.01 5.50 0.03 -0.15 -0.07 -0.04 0.03 -0.45 0.21 0.05 0.16 -0.04 -0.09 -0.11 3.36 -0.09 -2.18 -0.24 -0.37 -1.90
CGPT1.3 18.06 100.00 -2.22 -0.29 0.08 7.46 0.14 -0.30 0.10 -0.18 0.74 -0.79 0.07 0.23 0.23 0.04 0.02 0.00 5.10 -0.08 -1.34 0.03 -0.18 -1.17
Neo2.7 -1.07 -2.22 100.00 0.07 0.11 -1.13 0.08 0.04 -0.12 0.06 0.41 0.00 0.13 0.06 -0.00 0.10 0.09 -0.04 -0.77 0.08 -0.59 -0.15 0.00 -0.26
CGLM -0.67 -0.29 0.07 100.00 0.18 -1.07 -0.01 0.15 0.14 -0.12 0.04 0.37 -0.18 -0.09 -0.14 0.10 -0.01 0.08 -0.73 -0.07 0.09 0.18 -1.04 -0.40
CGLM2 0.01 0.08 0.11 0.18 100.00 -0.05 0.75 0.11 0.14 0.10 0.92 0.79 0.24 0.11 0.11 0.69 0.09 0.55 -0.07 -0.02 -0.12 -0.03 -0.01 -0.03
OPT6.7 5.50 7.46 -1.13 -1.07 -0.05 100.00 0.45 -0.36 -0.14 0.02 1.31 -1.09 -0.06 -0.06 0.41 0.17 -0.23 0.09 46.29 -0.38 -2.66 -0.03 -0.17 -1.34
Pythia6.9 0.03 0.14 0.08 -0.01 0.75 0.45 100.00 -0.00 0.01 0.02 2.37 0.55 -0.06 0.04 0.01 1.58 -0.02 0.88 0.29 -0.00 -0.03 -0.01 -0.01 -0.13
LM7 -0.15 -0.30 0.04 0.15 0.11 -0.36 -0.00 100.00 0.60 3.16 0.06 0.08 0.03 0.32 0.32 0.02 2.07 0.00 0.04 1.15 0.19 1.59 0.06 0.06
Qwen -0.07 0.10 -0.12 0.14 0.14 -0.14 0.01 0.60 100.00 0.53 -0.02 0.01 0.48 0.32 0.41 0.04 0.42 0.03 -0.12 -0.20 0.21 0.03 0.11 0.11
LM27 -0.04 -0.18 0.06 -0.12 0.10 0.02 0.02 3.16 0.53 100.00 -0.04 -0.09 -0.13 0.39 0.35 -0.03 1.67 0.02 0.37 1.77 0.55 1.71 0.15 0.03
RedPaj 0.03 0.74 0.41 0.04 0.92 1.31 2.37 0.06 -0.02 -0.04 100.00 0.35 0.02 0.06 0.09 2.08 0.03 1.24 0.68 -0.13 0.02 0.03 0.21 0.02
Bloom -0.45 -0.79 0.00 0.37 0.79 -1.09 0.55 0.08 0.01 -0.09 0.35 100.00 -0.01 0.09 -0.13 0.48 0.07 0.26 -0.68 0.02 -0.42 -0.08 -0.00 -0.05
Internlm 0.21 0.07 0.13 -0.18 0.24 -0.06 -0.06 0.03 0.48 -0.13 0.02 -0.01 100.00 0.18 0.21 0.02 0.08 -0.03 -0.64 -0.31 -0.13 -0.29 0.00 0.05
OLM 0.05 0.23 0.06 -0.09 0.11 -0.06 0.04 0.32 0.32 0.39 0.06 0.09 0.18 100.00 0.21 0.03 0.27 0.02 0.01 0.19 0.02 0.06 0.08 -0.01
Baichuan 0.16 0.23 -0.00 -0.14 0.11 0.41 0.01 0.32 0.41 0.35 0.09 -0.13 0.21 0.21 100.00 0.00 0.28 0.02 0.21 0.10 -0.25 0.10 -0.16 0.09
Pythia12 -0.04 0.04 0.10 0.10 0.69 0.17 1.58 0.02 0.04 -0.03 2.08 0.48 0.02 0.03 0.00 100.00 -0.02 0.78 0.08 -0.02 0.01 -0.04 -0.03 -0.09
LM13 -0.09 0.02 0.09 -0.01 0.09 -0.23 -0.02 2.07 0.42 1.67 0.03 0.07 0.08 0.27 0.28 -0.02 100.00 -0.01 0.13 0.39 0.21 0.88 0.07 -0.08
Neox20 -0.11 0.00 -0.04 0.08 0.55 0.09 0.88 0.00 0.03 0.02 1.24 0.26 -0.03 0.02 0.02 0.78 -0.01 100.00 0.04 -0.03 -0.08 0.02 -0.16 0.01
OPT30 3.36 5.10 -0.77 -0.73 -0.07 46.29 0.29 0.04 -0.12 0.37 0.68 -0.68 -0.64 0.01 0.21 0.08 0.13 0.04 100.00 0.12 -1.29 0.56 -0.06 -1.16
LM30 -0.09 -0.08 0.08 -0.07 -0.02 -0.38 -0.00 1.15 -0.20 1.77 -0.13 0.02 -0.31 0.19 0.10 -0.02 0.39 -0.03 0.12 100.00 0.11 2.45 -0.13 -0.12
Gala30 -2.18 -1.34 -0.59 0.09 -0.12 -2.66 -0.03 0.19 0.21 0.55 0.02 -0.42 -0.13 0.02 -0.25 0.01 0.21 -0.08 -1.29 0.11 100.00 -0.03 18.49 0.74
LM65 -0.24 0.03 -0.15 0.18 -0.03 -0.03 -0.01 1.59 0.03 1.71 0.03 -0.08 -0.29 0.06 0.10 -0.04 0.88 0.02 0.56 2.45 -0.03 100.00 -0.13 -0.01
Gala120 -0.37 -0.18 0.00 -1.04 -0.01 -0.17 -0.01 0.06 0.11 0.15 0.21 -0.00 0.00 0.08 -0.16 -0.03 0.07 -0.16 -0.06 -0.13 18.49 -0.13 100.00 -0.09
Fal180 -1.90 -1.17 -0.26 -0.40 -0.03 -1.34 -0.13 0.06 0.11 0.03 0.02 -0.05 0.05 -0.01 0.09 -0.09 -0.08 0.01 -1.16 -0.12 0.74 -0.01 -0.09 100.00
Table 4: ICS between 24 open-sourced LLMs with sizes ranging from 774M to 180B. Abbreviations used include ”GPT2” for ”GPT2-Large,” ”CGPT1.3” for ”Cerebras-GPT-1.3B,” ”Neo2.7” for ”GPT-Neo-2.7B,” ”CGLM” for ”ChatGLM-6B,” ”CGLM2” for ”ChatGLM2-6B,” ”OPT6.7” for ”OPT-6.7B,” ”Pythia6.9” for ”Pythia-6.9B,” ”LM7” for ”LLaMA-7B,” ”Qwen” for ”Qwen-7B,” ”LM27” for ”LLaMA2-7B2,” ”RedPaj” for ”RedPajama-7B,” ”Bloom” for ”Bloom-7B,” ”Internlm” for ”InternLM-7B,” ”OLM” for ”OpenLLaMA-7B,” ”Baichuan” for ”Baichuan-7B,” ”Pythia12” for ”Pythia-12B,” ”LM13” for ”LLaMA-13B,” ”Neox20” for ”GPT-NeoX-20B,” ”OPT30” for ”OPT-30B,” ”LM30” for ”LLaMA-30B,” ”Gala30” for ”Galactica-30B,” ”LM65” for ”LLaMA-65B,” ”Gala120” for ”Galactica-120B,” and ”Fal180” for ”Falcon-180B” respectively. From left to right LLMs are sorted by parameter size from smallest to largest.
Figure 7: Fingerprints of 24 independently trained LLMs.
5.4 Independently Trained LLMs in Smaller Scale
To validate the uniqueness and stability of the parameter direction of LLMs trained from scratch, we independently trained GPT-NeoX-350M (Black et al., 2022) models on a subset of the Pile dataset (Gao et al., 2020). First, we examined whether different parameter initializations merely caused by global random seeds result in distinct parameter directions. Second, we explored the variation in the model’s parameter vector direction during pretraining.
5.4.1 GPT-NeoX Models with Different Global Seeds
ICS Seed=1 Seed=2 Seed=3 Seed=4 Seed=1 100.00 2.08 2.23 2.08 Seed=2 2.08 100.00 2.40 2.26 Seed=3 2.23 2.40 100.00 2.29 Seed=4 2.08 2.26 2.29 100.00 Table 5: ICS values between GPT-NeoX models with different global seeds
[Uncaptioned image] Figure 8: Fingerprints of GPT-NeoX models trained with varying global seeds
5.4.2 Model’s Parameter Vector Direction’s Variations During Pretraining
Cosine similarity 56.23 84.65 88.95 90.96 92.13 93.22 94.25

### v3

@@ 4 Mapping the Invariant Terms to Image and Publish it through ZKP
4.1 Training
The encoder is the only component that needs to be trained in our fingerprinting model. Note that we don’t need to use any real LLM weights for training the encoder (all the training data are synthesized by randomly sampled matrices), as it only needs to learn a locality-preserving mapping between the input tensor and the output Gaussian vector. We adopt contrastive learning to learn locality-preserving mapping. To render the output vector to be Gaussian, we adopt the standard GAN (Karras et al., 2019) training scheme. (c.f. Appendix B for details of data synthesis and the whole training process.)
@@ 4.2 Inference
In the inference stage, the encoder takes the invariant terms from real LLMs and outputs ${\bm{v}}$. One image generator converts ${\bm{v}}$ into a natural image. In principle, any image generator that takes a Gaussian input and has the locality-preserving property would fit here. By visually representing the invariant terms as fingerprints, we can easily identify base models based on their fingerprint images, enabling reliable tracking of model origins. In this paper, we employ the StyleGAN2 generator pretrained on the AFHQ (Choi et al., 2020) dog dataset to generate natural images, we detailed it in Appendix E.
4.3 Zero-knowledge Proof for Fingerprints
In our black-box setting, users are unable to access the model parameters, which presents a significant challenge in ensuring the fingerprint is genuinely derived from the claimed LLM parameters. To address this issue, we employ zero-knowledge proof, a cryptographic technique that allows the prover to convince the verifier that a statement is true without revealing any information beyond the statement’s validity (Ben-Sasson et al., 2013; Goldwasser et al., 2019; Chiesa et al., 2020).
The manufacturer generates a publicly verifiable zero-knowledge proof along with computing the fingerprint, ensuring two critical aspects: (1) the input parameters indeed originate from the specific LLM the manufacturer claims, thereby safeguarding against substitution attack. (c.f. Appendix D for detailed discussion of substitution attack.) (2) the human-readable fingerprint is calculated correctly, confirming the fingerprint is genuinely derived from the LLM parameters. The detailed zero-knowledge proof generation process is as follows:
  1. 1.
    Select a random number $t$, commit to LLM parameters (which we denote by $model$) and input $\hat{{\bm{X}}}$, $\operatorname{commit}(model,\hat{{\bm{X}}},t)=\mathbf{cm}$. The commitment $\mathbf{cm}$ is public and does not reveal any information about the model.
  2. 2.
    While calculating fingerprint, generate a ZK proof $\pi_{1}$ prove that the manufacturer knows $model,\hat{{\bm{X}}},t$ s.t.
    1. (a)
      $model$ is the claimed LLM parameters and $\hat{{\bm{X}}}$ satisfy $\operatorname{commit}(model,\hat{{\bm{X}}},t)=\mathbf{cm}$;
    2. (b)
      The last two layers parameters ${\bm{W}}_{Q},{\bm{W}}_{K},{\bm{W}}_{V},{\bm{W}}_{O},{\bm{W}}_{1},{\bm{W}}_{2}$ in $model$ and input $\hat{{\bm{X}}}$ satisfy
    3. (c)
      The output human-readable fingerprint is indeed calculated from the invariant terms above.
As above( item 1. and item 2.a), the manufacturers commit to the claimed model and publish the commitment first, which is a conventional approach to ensure the parameters are not altered during the proof generation. All subsequent proof and inference processes will be carried out with this commitment, and anyone can verify if the model parameters used match those sealed within the commitment. The steps item 2.b and item 2.c are to ensure that the invariant terms and fingerprint are correctly calculated. Anyone who gets proof $\pi_{1}$ and the commitment $\mathbf{cm}$ can verify that the fingerprint is calculated based on LLM.
Moreover, we also provide a limited quantitative comparison scheme, which supports one-to-one comparison with open-source models. The manufacturers calculate the cosine similarity of invariant terms and give the zero-knowledge proof $\pi_{2}$ of this calculation process. Anyone who gets the proof $\pi_{2}$ and the open-source model can verify the cosine similarity without learning the private model.
@@ 5 Experiments
Our experiment is twofold. First, we validated the effectiveness and robustness of invariant terms for identifying the base model. Second, we generated fingerprints based on invariant terms for 80 LLMs and quantitatively assessed their discrimination ability through a human subject study.
5.1 Effectiveness and Robustness of Invariant Terms
In this subsection, we validate the effectiveness and robustness of invariant terms in identifying base model through four key experiments. First, we compute the Invariant Terms’ Cosine Similarity (ICS) between 8 widely used open-sourced LLM base models and their offspring models (including heavily continue-pretrained models), verifying its robustness against subsequent training processes. Second, we conduct extensive experiments on additional open-sourced LLMs, showcasing low ICS between 28 independently trained models. Third, we gather 51 offspring models and calculate the accuracy of correctly identifying the base model. Finally, we compare our methods with two latest baselines.
5.1.1 High ICS between Base LLMs and Their Offspring Models
ICS Falcon-40B LLaMA2-13B MPT-30B LLaMA2-7B Qwen-7B Baichuan-13B InternLM-7B
Offspring1 99.61 99.50 99.99 99.47 98.98 99.76 99.28
Offspring2 99.69 99.49 99.99 99.41 99.71 99.98 99.02
Table 2: The ICS between offspring models and their corresponding base model.
First, we perform experiments on 7 commonly used open-sourced LLMs, ranging in size from 7B to 40B. The 7 base models considered are Falcon-40B (Almazrouei et al., 2023), MPT-30B (Lin et al., 2022), LLaMA2-7B, 13B, Qwen-7B (Bai et al., 2023), Internlm-7B and Baichuan-13B. For each of these base models, we collect 2 popular offspring models. We extract the invariant terms for all these models and calculate the ICS for each offspring model w.r.t. its base model (Table 2). Remarkably, all offspring models exhibit very high ICS, with an average ICS of $\mathbf{99.56}$.
Second, we leverage the LLaMA-7B base model as a testing ground to assess the robustness of invariant terms under diverse subsequent training processes. We include 10 offspring models detailed in Section 3.1.1 and add Beaver, Guanaco (Dettmers et al., 2023), and BiLLa (Li, 2023) to the collection. See Appendix Table 10 for detailed descriptions. We extract invariant terms following the previous settings and compute the cosine similarity of the invariant terms (ICS) between each pair of models. Despite undergoing various training paradigms, such as RLHF, SFT, modality extension, and continued pretraining in a new language, we observe a high degree of similarity (Table 3), with an average ICS of $\mathbf{94.14}$.
@@ 5.1.1 High ICS between Base LLMs and Their Offspring Models
Table 3: The cosine similarities of invariant terms (ICS) between various pairs of LLaMA-7B and its offspring models. Abbreviations: MedAlpaca (MAlpaca), Alpaca-Lora (AlpacaL), MiniGPT-4 (MiGPT), WizardLM (Wizard), Chinese-Alpaca (CAlpaca), Chinese-LLaMA (CLLaMA).
5.1.2 Low ICS between 28 Independently Trained LLMs
Besides the aforementioned base models, we assemble a comprehensive collection of 28 open-sourced LLMs, ranging in size from 774M (GPT2-Large) to 180B (Falcon-180B). Please refer to Appendix F for details. We extract invariant terms and calculate ICS between each pair of models. Notably, the similarities between different models were consistently low, with an average ICS of $\mathbf{0.38}$, affirming the effectiveness of invariant terms. (c.f. Appendix Table 7 for detailed ICSs.)
5.1.3 Accuracy in Identify 51 Offspring Models’ Base Model
To assess the effectiveness of our method, we gathered 51 offspring models derived from 18 distinct base models. (c.f.  Table 10 for detailed list and description.) Calculating the ICS between each offspring model and the 18 base models, we predicted the base model with the highest ICS. Comparing these predictions with the ground truth, our method accurately identified the base models for all 51 offspring LLMs, achieving $\mathbf{100\%}$ accuracy.
FSR MiGPT Alpaca MAlpaca Vicuna Wizard Baize AlpacaL CAlpaca Koala CLLaMA Beaver Guanaco BiLLa Avg.
Trap 0 8.04 24.14 2.30 17.24 44.83 39.08 1.15 0 0 8.05 10.34 0 11.94
[…]
Ours 100 100 100 100 100 100 100 100 100 100 100 100 100 100.00
[…]
Figure 5: Fingerprints of 7 different base models (in the first row) and their corresponding offspring models (the lower two rows) are presented. The base model’s name is omitted in the offspring models.
5.1.4 Comparing to Latest Fingerprinting Methods
[…]
5.2 Discrimination Ability of Human-readable Fingerprints
Based on previous invariant terms, we employ the fingerprinting model illustrated in Section 4 to generate and publish human-readable fingerprints for previously mentioned LLMs. In Figure 5, there are fingerprints of the 7 independently trained LLMs and their offspring models (Section 5.1.1). Notably, for all the offspring models, their fingerprints closely resemble those of their base models. On the other hand, LLMs based on different models yield highly distinctive fingerprints, encompassing various appearances and breeds of dogs. Due to space limit, the fingerprints of LLaMA family models, the rest offspring models, and the 28 independently trained LLMs are listed in Appendix H.
Furthermore, we conducted a human subject study and yielded a $\mathbf{94.74\%}$ accuracy rate (c.f. Appendix I for details), quantitatively demonstrate the discrimination ability of our generated fingerprints. Although using human-readable fingerprints introduces minor losses, manufacturers can provide one-to-one comparison results with proof to make up for this loss of misjudgment.
Except for the aforementioned experiments, we independently train LLMs on a smaller scale to provide further validation for our method. (c.f. Appendix G)